OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x - y = 3

OpenStudy (bella2112):

4x + 2y = 12 x - y = 3 then it might equal 12-9=3 try it

OpenStudy (bella2112):

if im wrong srry

OpenStudy (anonymous):

that is way wrong lol

OpenStudy (bella2112):

well srry

OpenStudy (bella2112):

im not in your grade so yeah

OpenStudy (anonymous):

what grade are you in then

OpenStudy (anonymous):

getting a little mad you can do freshman math XD

OpenStudy (bella2112):

im in 5th grade so yeah

OpenStudy (bella2112):

xD